### v1

@@ 6.4 Estimate of the radius $R$
On one hand, from the afterglow perspective the values for the density of the circumburst medium may play a significant role in the derivation of $\Gamma$ for the wind–like medium, as noted before, while on the other hand, from the pair–production argument perspective, we are dependent from the assumptions on the variability timescale. In deriving the bulk Lorentz factor reported in Sec. 5, we set $t_{\rm var}$ = 0.1 s for all GRBs, based on the average value of the observed distribution of variability timescale for a sample of Fermi  detected GRBs (Golkhou et al. 2015). However, this value may likely not correspond to the intrinsic value of the variability timescale for each burst in our sample and this may affect the estimated $\Gamma$ and, consequently, also the value of the distance $R$ of the GRB emitting region from the central engine. For every GRBs of our sample, the implied value of the radius $R$ (following the usual $R=2ct_{\rm var}\Gamma^{2}/(1+z)$) is reported in the first row of the fourth column of the Table 3.
The distance $R$ is of one least known and often assumed parameters of the GRB standard model. For those GRBs belonging to both samples, we can take the advantage of having the bulk Lorentz factor constrained from an independent method (e.g. the afterglow lightcurve) and constrain the radius $R$. Indeed, we can invert Eq. 3 and use the $\Gamma$ value inferred from the afterglow onset to derive the value (or a lower–limit) of the radius $R$, applying the compactness argument, without making assumptions on the variability timescale.
@@
This method has the advantage of being independent from the assumptions on (or the measurements of) the variability timescale $t_{\rm var}$. It is always possible, however, to compute the variability timescale a posteriori for a given $\Gamma$ and $R$: indeed, in the last column of Table 3 we report the variability timescales implied by the corresponding values of $\Gamma$ and $R$, when this method has been applied.
@@
If the GRB spectrum shows an exponential cutoff, then the cutoff energy $E_{\rm cutoff}$ allows to derive the estimate of $R$ (by setting $\tau_{\gamma\gamma}(E_{\rm cutoff})=1$, as before). If instead the spectrum is a single power–law above the peak, then the lower–limit on the cutoff energy will allow to derive only a lower–limit on $R$. We stress, however, that this method holds under the assumption that the bulk Lorentz factor $\Gamma$ of the jet derived from the afterglow onset corresponds to average value of the distribution of bulk Lorentz factors of the jet during the prompt emission.
@@
We show in Table 3 the relative estimate of $R$, for GRB 080916 and GRB 110731, following Eq. 8 and using the values of $\Gamma$ obtained for the homogeneous and for the wind–like medium by Ghirlanda et al. 2018. The two estimates of the radius derived in the two cases are quite different: $R$ in the wind–like case is greater by two orders of magnitude than in the homogeneous case. This is due to the corresponding $\Gamma$ values, which differ by a factor of $\sim$2–3. As shown in Eq. 8, $R\propto\Gamma^{-2-2\beta_{e}}$: for a typical energy index $\beta_{e}\sim 1.5$, we have $R\propto\Gamma^{-5}$. In both cases however, the values of the distance $R$ are found to be $R\gtrsim 10^{13-15}$ cm. This might be suggesting that, although still consistent with it, the distance of the emitting region from the central engine could be slightly larger than the radius typically assumed in the GRB standard model for the internal shocks to occur. An additional theoretical lower–limit on the distance $R$ to take into account can be derived from the photospheric radius $R_{\rm ph}$ (Daigne & Mochkovitch 2002), below which the emitting region is opaque. Given the isotropic luminosity of each burst and assuming a typical prompt efficiency of $\eta=0.2$ (Nava et al. 2014; Beniamini et al. 2015), we checked that the values of $R$ derived using the pair–production argument for each bulk Lorentz factor considered are safely above the associated values of the photospheric radius (in the range $R_{\rm ph}\sim 10^{11}-10^{12}$ cm).
@@
On the other side, $R$ is limited also downward by the variability timescale of the prompt lightcurve. Assuming the standard scenario where the emission is produced by expanding spherical shells, the short (order of $\sim$ tens or hundreds of ms, see e.g. Golkhou et al. (2015)) variability timescales observed in the prompt emission constrains the possible combinations of bulk Lorentz factor $\Gamma$ and distance $R$ from the central engine (i.e. $t_{\rm var}=R/2c\Gamma^{2}\sim 0.167\,R_{14}/2c\Gamma_{2}^{2}$ s). If the prompt emission site is moved at larger distances from the central engine, e.g. $R\sim 10^{16}$ cm, it becomes hard to obtain the observed fast variability timescales without invoking extreme bulk Lorentz factors (i.e. $\Gamma\sim 1000$).
@@
Table 3: Bulk Lorentz factors $\Gamma$ for each GRB co-detected by GBM and LAT analyzed in this work, applying the compactness argument. Depending on the presence or absence of a cutoff in the high-energy spectrum, these values are measurements or lower–limits, respectively. For each GRB, the table list the name, the method applied to derive $\Gamma$, the corresponding value of bulk Lorentz factor $\Gamma$, the implied value of the distance R from the central engine and the variability timescale, which we assumed to be 100 ms for each GRB. For the only two GRB in our sample (GRB 080916C and GRB 110131) with also an independent estimate of $\Gamma$ derived from the afterglow onset in Ghirlanda et al. 2018, we reported those $\Gamma$ values for the ISM and wind–like cases and the implied lower–limits on the distance R when combined with the pair–production opacity argument (following Eq.8). In those two cases, we also report the different variability timescale implied by the corresponding values of R and $\Gamma$, following $t_{\rm var}=R(1+z)/(2c\Gamma^{2})$.
@@
Figure 8: Constraints on the bulk Lorentz factor $\Gamma$ and distance $R$ of the emitting region from the central engine, for the 22 bright GRBs analyzed in this work. The measured values (blue crosses) or lower limits (red arrows) on $\Gamma$, and consequently on $R$ (assuming $t_{\rm var}=100$ ms), are derived according to the presence or absence of a high-energy cutoff in the spectrum, respectively. For each GRB with a measured value of $\Gamma$, the green stripes represent the region of the parameter space allowed by the compactness argument for any given $\Gamma$, using only the prompt spectral properties and following Eq. 8. When available, the combination of the information from the prompt and from the afterglow allows to break the degeneracy and better constrain the distance $R$. The four black arrows represent the cases of GRB 080916 and GRB 110731, where an independent estimate of $\Gamma$ from the afterglow onset was taken from Ghirlanda et al. 2018, allowing to put constraints on the distance $R$. Corresponding lower limits on variability timescales are provided for each case. Each error bar represents the 1-$\sigma$ error on measurements.
In our analysis, this method for the estimate of the distance $R$ has been applied to only two GRBs of our sample which satisfied the criterion of having $\Gamma$ estimated in an independent way. However, Eq. 8 is general and can be applied to any other GRB which has the bulk Lorentz factor $\Gamma$ inferred in a independent way and whose spectral parameters of the prompt emission are known. Indeed, this method has been recently applied in Mei et al. (2022) to GRB 220101A, a bright burst ($E_{\rm iso}\sim 3\times 10^{54}$ erg) detected by Fermi/GBM and Fermi/LAT. The joint multi–wavelength analysis of the prompt emission spectrum allowed to characterize the full spectral shape from X–rays to GeV energies, revealing the detection of a spectral cutoff at $E_{\rm cutoff}=80^{+44}_{-14}$ MeV (similar results found also in Scotton et al. 2023). Combining the information gathered from the prompt emission spectrum and the afterglow lightcurve, the authors derived stringent constraints on the GRB prompt emission site, finding $\Gamma\sim 900$ and $R\sim 4.5\times 10^{13}$ cm (see their Figure 4). We note that this GRB did not belong to the Fermi/LAT Catalog at the time of the selection of the sample for this work. However, while it would have satisfied the criterion set on the fluence detected in GBM (its fluence being $F=6.03\pm 0.02\times 10^{-5}\,\,\rm erg/cm^{2}$), its significance in LLE data is only $\sigma=14$, so it would not be part of our selected sample of bright LLE-detected bursts. Nonetheless, this burst clearly showed evidence of a cutoff at high–energies, implying that there might be other GRBs candidates with a cutoff in their prompt spectrum outside our S/N-selected sample.
@@ 7 Conclusions
In this work, we present the spectral analysis of the 22 brightest GRBs simultaneously detected by both instruments on board Fermi, GBM and LAT, covering a wide energy range, from 10 keV to 10 GeV.
@@
The joint analysis of GBM and LAT data for the 22 time-integrated prompt emission GRB spectra revealed that in 9 burst the spectral data significantly deviate from the extrapolation of the high–energy power–law, requiring the presence of an exponential cutoff at high energies, between $\sim$14 and 298 MeV. These results are consistent with the previous detections of exponential cutoffs in GRB prompt spectra reported in a few cases in the literature (e.g. Vianello et al. 2018; Tang et al. 2015). In the remaining 13 out of 22 GRBs of the sample, there is no significant evidence of a spectral cutoff in the data. Since the poor signal in the highest part of the spectrum might hamper the evidence of a cutoff, for each of the 13 GRBs we derived a lower–limit on the cutoff position.
@@
Interpreting the exponential cutoff observed in 9 out of 22 GRBs as a sign of the pair-production opacity, we derive the estimate (and the lower–limit for the remaining 13 GRBs not showing the cutoff) of the jet bulk Lorentz factor $\Gamma$ during the prompt phase (see Fig. 5). The distribution of the estimates (lower–limits) of $\Gamma$ inferred from the pair-production argument spans the range of values 130–333 (136–745), which are compatible with the typical values assumed in the GRB standard model and are consistent with those derived from the same opacity argument reported in the literature.
@@
By comparing the $\Gamma$ values inferred from the opacity argument reported in this work with the ones inferred from the afterglow onset time reported in Ghirlanda et al. 2018, we found that our results are consistent with the distribution of $\Gamma$ derived for the homogeneous medium case (see Fig. 7). In the wind case, instead, the distribution of bulk Lorentz factors extends toward lower values of $\Gamma$, which would imply lower values of the cutoff energy, and consequently a fainter flux in the LAT range. Therefore, the absence of such lower values in the $\Gamma$ distribution of LAT-detected bursts could be due to a selection effect, hampering the consistency with the bulk Lorentz factors derived in the wind case. The distribution of $\Gamma$ may also extend towards higher values, as indeed suggested by the lower-limits on $\Gamma$ values, but their observation is hampered by the limited energy range of the current instruments.
@@
Combining all the spectral fits performed in this analysis, we derived the distribution of the slope $\beta$ of the high-energy power–law, with median value $\langle\beta\rangle$ = -2.39 ($\sigma=0.56$). By comparing it with the distribution of $\beta$ from the Fermi/GBM Catalog, we found steeper $\beta$ values with respect to the ones reported in the Catalog. This could be an indication that the true slope of the GRBs prompt emission spectrum is softer than what is usually found, and it is likely due to the addition of the LLE and LAT data in our analysis, that allow to better constrain the high–energy power–law slope of spectrum.
@@
From the slopes $\beta$ of the power–law modelling the high–energy spectrum of the 22 GRBs analyzed, we derived the corresponding slopes $p$ of the underlying non-thermal distribution of accelerated particles, assuming synchrotron emission in the fast-cooling regime. The distribution of $p$ is quite broad and does not cluster around a universal value. We found that the distribution of $p$ has a median value of $p=2.79$ with a tail reaching values of $p\sim$ 5–7. Steep values of $p$, coupled also with the large values of $\gamma_{\rm min}$ found from independent synchrotron modeling of other bursts in the literature (e.g. Ronchi et al. 2020; Ryde et al. 2022), require an efficient acceleration mechanism (mildly relativistic shocks or magnetic reconnection) giving rise to a steep energy distribution of accelerated particles. Given the theoretical uncertainties on the energy distribution of accelerated particles in mildly-relativistic shocks, these results provide useful observational benchmarks for the development of the theory of particle acceleration applied to the prompt emission case.
@@
Our results show how crucial is the inclusion, when possible, of the LLE (30–100 MeV) and LAT data (¿100 MeV) to the study of a GRB prompt spectrum detected by the GBM (typically sampled in the energy range 10 keV–10 MeV). A broadband study of the prompt emission spectrum can reveal significant deviations from the standard spectral shape often found when analyzing the spectra in a narrower energy range and with simpler functions (see the solid blue line in Fig. 2). As we show in this work, once obtained the crucial information from the high-energy prompt emission data, one can exploit their full potential by combining them with the information derived from the onset of the afterglow component, which allows to derive the location of the emitting region, without using the additional assumption on the variability timescale. In the only two GRBs of our sample satisfying the criteria, we used the independent estimates of $\Gamma$ derived from the afterglow onset reported in Ghirlanda et al. 2018. Depending on the assumptions on the homogeneous or wind medium cases, we found values of $R>10^{13}-10^{15}$ cm, which are compatibles with the location of the prompt emission region predicted by the GRB standard model.
@@
This method provides a powerful tool to restrict the range of distance $R$ of the prompt emitting region from the central engine, a key parameter in the GRB standard model still suffering from large uncertainties. Given the encouraging results presented in this work (and also applied in Mei et al. (2022)), we plan to apply this method to other GRBs that satisfy the criteria, namely whose prompt emission has been detected by LAT and with a measured afterglow onset time. This would significantly help in constraining the allowed region of the parameters space for prompt emission and distinguishing different dissipation models proposed in the literature, which invoke different emitting region distances and attainable bulk Lorentz factors. The investigation of these interesting issues will be the subject of a forthcoming publication.

### v2

@@ 6.4 Estimate of the radius $R$
On one hand, from the afterglow perspective the values for the density of the circumburst medium may play a significant role in the derivation of $\Gamma$ for the wind–like medium, as noted before, while on the other hand, from the pair–production argument perspective, we are considering the internal-shock framework and assuming a variability timescale. In deriving the bulk Lorentz factor reported in Sec. 5, we set $t_{\rm var}$ = 0.1 s for all GRBs, based on the average value of the observed distribution of variability timescale for a sample of Fermi  detected GRBs (Golkhou et al. 2015). For every GRBs of our sample, the implied value of the radius $R$ (following the usual $R=2ct_{\rm var}\Gamma^{2}/(1+z)$) is reported in the first row of the fourth column of the Table 3.
However, this value may likely not correspond to the intrinsic value of the variability timescale for each burst in our sample and this may affect the estimated $\Gamma$ and, consequently, also the value of the distance $R$ of the GRB emitting region from the central engine. The distance $R$ is one of the least known and often assumed (via the variability timescale) parameters of the GRB standard model. Indeed, as pointed out in Gupta & Zhang 2008 and used in for example Zhang & Pe’er 2009 and Chand et al. 2020, a thorough treatment free from the assumptions on the variability timescale and on the internal shocks would be to build a two-dimensional constraint in the $\Gamma-R$ plane. To break the degeneracy, we can then take advantage of having the bulk Lorentz factor constrained from an independent method (e.g., the afterglow light curve), when available, and constrain the radius $R$.
Indeed, one can invert Eq. 3 (see also Eq.(13) in Gupta & Zhang 2008) and use the $\Gamma$ value inferred from the afterglow onset to derive the value (or a lower–limit) of the radius $R$, applying the compactness argument, without making assumptions on the variability timescale.
@@
This method has the advantage of being independent of the assumptions on (or the measurements of) the variability timescale $t_{\rm var}$. It is always possible, however, to compute the variability timescale a posteriori for a given $\Gamma$ and $R$: indeed, in the last column of Table 3 we report the variability timescales implied by the corresponding values of $\Gamma$ and $R$, when this method has been applied.
@@
If the GRB spectrum shows an exponential cutoff, then the cutoff energy $E_{\rm cutoff}$ allows to derive the estimate of $R$ (by setting $\tau_{\gamma\gamma}(E_{\rm cutoff})=1$, as before). If instead, the spectrum is a single power–law above the peak, then the lower–limit on the cutoff energy will allow us to derive only a lower–limit on $R$. We stress, however, that this method holds under the assumption that the bulk Lorentz factor $\Gamma$ of the jet derived from the afterglow onset corresponds to the average value of the distribution of bulk Lorentz factors of the jet during the prompt emission. If instead, the jet would have a $\Gamma$ at the deceleration higher than the average value of $\Gamma$ of the prompt emission (e.g., due to magnetic acceleration, as suggested in Poynting-flux-dominated models), then this assumption would not hold.
@@
We show in Table 3 the relative estimate of $R$, for GRB 080916 and GRB 110731, following Eq. 8 and using the values of $\Gamma$ obtained for the homogeneous and for the wind–like medium by Ghirlanda et al. 2018. The two estimates of the radius derived in the two cases are quite different: $R$ in the wind–like case is greater by two orders of magnitude than in the homogeneous case. This is due to the corresponding $\Gamma$ values, which differ by a factor of $\sim$2–3. As shown in Eq. 8, $R\propto\Gamma^{-2-2\beta_{e}}$: for a typical energy index $\beta_{e}\sim 1.5$, we have $R\propto\Gamma^{-5}$. In both cases, however, the values of the distance $R$ are found to be $R\gtrsim 10^{13-15}$ cm. This might suggest that, although still consistent with it, the distance of the emitting region from the central engine could be slightly larger than the radius typically assumed in the GRB standard model for the internal shocks to occur. We note that large radii ($R\sim 10^{15}$ cm) would be favored in Poynting-flux dominated models, such as the Internal-Collision-induced Magnetic Reconnection and Turbulence (ICMART) model (Zhang & Yan 2011), in order to reach the condition of triggering a reconnection/turbulence cascade. An additional theoretical lower–limit on the distance $R$ to take into account can be derived from the photospheric radius $R_{\rm ph}$ (Daigne & Mochkovitch 2002), below which the emitting region is opaque. Given the isotropic luminosity of each burst and assuming a typical prompt efficiency of $\eta=0.2$ (Nava et al. 2014; Beniamini et al. 2015), we checked that the values of $R$ derived using the pair–production argument for each bulk Lorentz factor obtained are safely above the associated values of the photospheric radius (in the range $R_{\rm ph}\sim 1.6\times 10^{12}-6.8\times 10^{13}$ cm).
@@
On the other side, $R$ is limited also downward by the variability timescale of the prompt light curve. Assuming the standard scenario where the emission is produced by expanding spherical shells, the short (order of approximately tens or hundreds of ms, see e.g., Golkhou et al. (2015)) variability timescales observed in the prompt emission constrain the possible combinations of bulk Lorentz factor $\Gamma$ and distance $R$ from the central engine (i.e. $t_{\rm var}=R/2c\Gamma^{2}\sim 0.167\,R_{14}/2c\Gamma_{2}^{2}$ s). If the prompt emission site is moved at larger distances from the central engine, for example $R\sim 10^{16}$ cm, it becomes hard to obtain the observed fast variability timescales without invoking extreme bulk Lorentz factors (i.e. $\Gamma\sim 1000$). A somewhat similar way out has been proposed in magnetic dissipation models (e.g., ICMART, Zhang et al. 2011), where mini-jets moving at $\Gamma_{\rm mini-jets}$ are used to produce the short observed variability while keeping the emitting region at large radii ($R\sim 10^{15-16}$ cm). Nevertheless, the recent findings on the low-energy breaks (Oganesyan et al. 2017, 2018; Ravasio et al. 2018, 2019) require a small magnetic field ($B^{\prime}\sim$1-10 G, in the leptonic scenario of the synchrotron interpretation) in the emitting region, and the idea of a magnetically dominated jet appears difficult to be reconciled with such low values (for other constraints on GRB prompt emission models based on Poynting-flux-dominated jets, see also Beniamini & Piran 2014).
@@
Table 3: Bulk Lorentz factors $\Gamma$ for each GRB that was co-detected by GBM and LAT and analyzed in this work, applying the compactness argument. Depending on the presence or absence of a cutoff in the high-energy spectrum, these values are measurements or lower–limits, respectively. For each GRB, the table list the name, the method applied to derive $\Gamma$, the corresponding value of bulk Lorentz factor $\Gamma$, the implied value of the distance R from the central engine and the variability timescale, which we assumed to be 100 ms for each GRB. For the only two GRB in our sample (GRB 080916C and GRB 110131) with also an independent estimate of $\Gamma$ derived from the afterglow onset in Ghirlanda et al. 2018, we reported those $\Gamma$ values for the ISM and wind–like cases and the implied lower–limits on the distance R when combined with the pair–production opacity argument (following Eq.8). In those two cases, we also report the different variability timescale implied by the corresponding values of R and $\Gamma$, following $t_{\rm var}=R(1+z)/(2c\Gamma^{2})$.
@@
Figure 8: Constraints on the bulk Lorentz factor $\Gamma$ and distance $R$ of the emitting region from the central engine, for the 22 bright GRBs analyzed in this work. The measured values (blue crosses) or lower limits (red arrows) on $\Gamma$, and consequently on $R$ (assuming $t_{\rm var}=100$ ms), are derived according to the presence or absence of a high-energy cutoff in the spectrum, respectively. For each GRB with a measured value of $\Gamma$, the green stripes represent the region of the parameter space allowed by the compactness argument for any given $\Gamma$, using only the prompt spectral properties and following Eq. 8. The combination of the information from the prompt and the afterglow, when available, allows us to break the degeneracy and better constrain the distance $R$. The four black arrows represent the cases of GRB 080916 and GRB 110731, where an independent estimate of $\Gamma$ from the afterglow onset was taken from Ghirlanda et al. 2018, allowing to put constraints on the distance $R$. Corresponding lower limits on variability timescales are provided for each case. Each error bar represents the 1-$\sigma$ error on measurements.
In our analysis, this method for the estimate of the distance $R$ has been applied to only two GRBs of our sample which satisfied the criterion of having $\Gamma$ estimated in an independent way. However, Eq. 8 is general and can be applied to any other GRB which has the bulk Lorentz factor $\Gamma$ inferred in a independent way and whose spectral parameters of the prompt emission are known. Indeed, this method has been recently applied in Mei et al. (2022) to GRB 220101A, a bright burst ($E_{\rm iso}\sim 3\times 10^{54}$ erg) detected by Fermi/GBM and Fermi/LAT. The joint multiwavelength analysis of the prompt emission spectrum allowed to characterize the full spectral shape from X–rays to GeV energies, revealing the detection of a spectral cutoff at $E_{\rm cutoff}=80^{+44}_{-14}$ MeV (similar results found also in Scotton et al. 2023). Combining the information gathered from the prompt emission spectrum and the afterglow light curve, the authors derived stringent constraints on the GRB prompt emission site, finding $\Gamma\sim 900$ and $R\sim 4.5\times 10^{13}$ cm (see their Figure 4). We note that this GRB did not belong to the Fermi/LAT catalog at the time of the selection of the sample for this work. However, while it would have satisfied the criterion set on the fluence detected in GBM (its fluence being $F=6.03\pm 0.02\times 10^{-5}\,\,\rm erg/cm^{2}$), its significance in LLE data is only $\sigma=14$, so it would not be part of our selected sample of bright LLE-detected bursts. Nonetheless, this burst clearly showed evidence of a cutoff at high–energies, implying that there might be other GRBs candidates with a cutoff in their prompt spectrum outside our S/N-selected sample.
@@ 7 Conclusions
In this work, we present the spectral analysis of the 22 brightest GRBs simultaneously detected by both instruments on board Fermi, GBM, and LAT, covering a wide energy range, from 10 keV to 10 GeV. The joint analysis of GBM and LAT data for the 22 time-integrated prompt emission GRB spectra revealed that in nine bursts the spectral data significantly deviate from the extrapolation of the high–energy power law, requiring the presence of an exponential cutoff at high energies, between $\sim$14 and 298 MeV. These results are consistent with the previous detections of exponential cutoffs in GRB prompt spectra reported in a few cases in the literature (e.g., Vianello et al. 2018; Tang et al. 2015). In the remaining 13 out of 22 GRBs of the sample, there is no significant evidence of a spectral cutoff in the data. Since the poor signal in the highest part of the spectrum might hamper the evidence of a cutoff, for each of the 13 GRBs we derived a lower limit on the cutoff position.
@@
Interpreting the exponential cutoff observed in nine out of 22 GRBs as a sign of the pair-production opacity, we derived the estimate (and the lower limit for the remaining 13 GRBs not showing the cutoff) of the jet bulk Lorentz factor $\Gamma$ during the prompt phase (see Fig. 5). The distribution of the estimates (lower limits) of $\Gamma$ inferred from the pair-production argument spans the range of values 130–333 (136–745), which are compatible with the typical values assumed in the GRB standard model and are consistent with those derived from the same opacity argument reported in the literature.
@@
By comparing the $\Gamma$ values inferred from the opacity argument reported in this work with the ones inferred from the afterglow onset time reported in Ghirlanda et al. 2018, we found that our results are consistent with the distribution of $\Gamma$ derived for the homogeneous medium case (see Fig. 7). In the wind case, instead, the distribution of bulk Lorentz factors extends toward lower values of $\Gamma$, which would imply lower values of the cutoff energy, and consequently a fainter flux in the LAT range. Therefore, the absence of such lower values in the $\Gamma$ distribution of LAT-detected bursts could be due to a selection effect hampering the consistency with the bulk Lorentz factors derived in the wind case. The distribution of $\Gamma$ may also extend toward higher values, as is indeed suggested by the lower limits on $\Gamma$ values, but their observation is hampered by the limited energy range of the current instruments.
@@
Combining all the spectral fits performed in this analysis, we derived the distribution of the slope $\beta$ of the high-energy power law, with median value $\langle\beta\rangle$ = -2.39 ($\sigma=0.56$). By comparing it with the distribution of $\beta$ from the Fermi/GBM catalog, we found steeper $\beta$ values with respect to the ones reported in the catalog. This could be an indication that the true slope of the GRBs’ prompt emission spectrum is softer than what is usually found, and it is likely due to the addition of the LLE and LAT data in our analysis, which allowed us to better constrain the high–energy power–law slope of the spectrum.
@@
From the slopes $\beta$ of the power law modeling the high–energy spectrum of the 22 GRBs analyzed, we derived the corresponding slopes $p$ of the underlying nonthermal distribution of accelerated particles, assuming synchrotron emission in the fast-cooling regime. The distribution of $p$ is quite broad and does not cluster around a universal value. We found that the distribution of $p$ has a median value of $p=2.79$ with a tail reaching values of $p\sim$ 5–7. Steep values of $p$, coupled also with the large values of $\gamma_{\rm min}$ found from independent synchrotron modeling of other bursts in the literature (e.g., Ronchi et al. 2020; Ryde et al. 2022), require an efficient acceleration mechanism (mildly relativistic shocks or magnetic reconnection) giving rise to a steep energy distribution of accelerated particles. Given the theoretical uncertainties on the energy distribution of accelerated particles in mildly relativistic shocks, these results provide useful observational benchmarks for the development of the theory of particle acceleration applied to the prompt emission case.
@@
Our results show how crucial the inclusion is, when possible, of the LLE (30–100 MeV) and LAT data (¿100 MeV) to the study of a GRB prompt spectrum detected by the GBM (typically sampled in the energy range 10 keV–10 MeV). A broadband study of the prompt emission spectrum can reveal significant deviations from the standard spectral shape often found when analyzing the spectra in a narrower energy range and with simpler functions (see the solid blue line in Fig. 2). As we show in this work, once the crucial information from the high-energy prompt emission data is obtained, one can exploit their full potential by combining them with the information derived from the onset of the afterglow component, which allows the location of the emitting region to be derived without using the additional assumption on the variability timescale (see also Gupta & Zhang 2008). In the only two GRBs of our sample satisfying the criteria, we used the independent estimates of $\Gamma$ derived from the afterglow onset reported in Ghirlanda et al. 2018. Depending on the assumptions on the homogeneous or wind medium cases, we found values of $R>10^{13}-10^{15}$ cm, which are compatible with the location of the prompt emission region predicted by the GRB standard model, although they may suggest larger distances, as expected in magnetic dissipation models (Lyutikov & Blandford 2003; Zhang et al. 2011) and also found in Zhang & Pe’er 2009 and Beniamini et al. 2018, for example.
@@
This method provides a powerful tool to restrict the range of distance $R$ of the prompt emitting region from the central engine, a key parameter in the GRB standard model still suffering from large uncertainties. Given the encouraging results presented in this work (and also applied in Mei et al. (2022)), we plan to apply this method to other GRBs that satisfy the criteria, namely whose prompt emission has been detected by LAT and with a measured afterglow onset time. This would significantly help in constraining the allowed region of the parameters’ space for prompt emission and distinguishing between different dissipation models proposed in the literature, which invoke different emitting region distances and attainable bulk Lorentz factors. The investigation of these interesting issues will be the subject of a forthcoming publication.